A basket contains some apples. Alice takes 1/2 of the apples and then places 15 of the apples back in the basket. Barry then takes 1/2 of the remaining apples and places 10 back in the basket.They find that each of them has the same number of apples. How many apples are left in the basket?
(A) 10 (B) 20 (C) 30 (D) 50 (E) Not enough information

Answer by richwmiller(17219)

(x/2-15)=(x/2+15)/2-10
b =(x/2+15)/2-10
a=x/2-15
k=x-a-b
We only need the first equation to solve for x the number of apples in the beginning
b= barry's apples at end
a= alice's apples at end
k equals number of apples left in basket at end
x= 50 apples in the beginning
a takes 25 and puts 15 back
a=10
basket has 40 now
b takes 20 and returns 10
b=10
basket has 30 now
Barry and Alice each have 10
30 left in the basket